OpenStudy (anonymous):

which of the following are solutions to the equation Check all that apply. 4x^2-12x+9=5

OpenStudy (anonymous):

a.\[x=\frac{- \sqrt{5} +3}{ 2 }\]

OpenStudy (anonymous):

B.\[x=\sqrt{4}-3\]

OpenStudy (anonymous):

C.\[x=\frac{ \sqrt{5}+3 }{ 2 }\]

OpenStudy (anonymous):

D.\[x=\sqrt{5}+\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

E.\[x=-\sqrt{5}+\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

F.\[x=-\sqrt{4}-3\]

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

can someone check my answer it the link

OpenStudy (anonymous):

\[4x^2-12x+9=5 \]\[(3-2 x)^2 = 5\]\[(2 x-3)^2 = 5\]Then take the square root!

OpenStudy (anonymous):

is the link trying to say \[x=\sqrt{11}+3\]

OpenStudy (ybarrap):

$$\tt \tt \begin{align*} \\ 4x^2-12x+9=5\\ 4x^2-12x+4=0\\ 4(x^2-3x+1)=0)\\ (x^2-3x+1)=0)\\ x=\frac{3\pm\sqrt{9-4}}{2}\\ x=\frac{3\pm\sqrt{5}}{2}\\ \end{align*}$$

OpenStudy (anonymous):

YUP! A and C

OpenStudy (anonymous):

and \[x=-\sqrt{11}+3\]

OpenStudy (anonymous):

No. It is not!

OpenStudy (anonymous):

Look at what @ybarrap said!